OpenStudy (anonymous):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)) y=(8/x^2)+4

OpenStudy (anonymous):

simple \[g(x)\] would take over the spot of the x in \[f(x)\] so if \[f(g(x))=y\] and \[y=(\frac{ 8 }{ x }^{2}) +4\] then \[g(x)=\frac{ 8 }{ x }^{2}\] and \[f(x)=x+4\]

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)) y= (7/x^2)+10 can you help me with too :)

OpenStudy (anonymous):

its the same concept in \[y=(\frac{ 7 }{ x }^{2})+10\]the stuff in the parentheses is what g(x) is equal to and f(x) is the same only replace the stuff in the parentheses for x